OpenStudy (anonymous):

bob used X sheets of a 500-sheet pack of typing paper. mike used twice as much as bob. pete used half as much as mike. which algebraic experssion represtents the amount of paper remaining A. 4X B. 3.5X C. 500-4X D. 600-3.5X PLEASE HELP

OpenStudy (anonymous):

D. 500-3.5X (MISTAKE)

OpenStudy (anonymous):

500 B = X M = 2B => M = 2X P = M/2 => P = 2X/2 = X ....ADDING ALL 500 - X - 2X - X 500 - 4X

OpenStudy (smurfy14):

ok nevermind the answer is C because m=2b p=1/2m therefore b+p=m

OpenStudy (anonymous):

There are 500 pieces of paper. Bob uses X amount and therefore at that point the amount remaining would be 500 minus that X amount or 500-X. If Mike uses twice as much as Bob then he is using 2X and therefore there is 500 - X - 2X. Pete then uses half of the amount that Mike used, which was 2X, meaning he used X. Leaving 500 - X - 2X -X. When when combine like terms we get 500 - 4X.